OpenStudy (anonymous):

Equation. Don't get it... 4(x-3)=22

OpenStudy (anonymous):

take 4 multiply by x and 3 4x-12=22 now just add 12 to both sides then divide by 4 then you will have x

OpenStudy (anonymous):

0.18

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4x=34 x=8.5

OpenStudy (anonymous):

I wrote down 4(x-3)=22 4x-12=22... N got lost from there ...

OpenStudy (anonymous):

ok 4x-12=22 +12 +12 you want to get x by itself and to do that u add 12 to both sides 4x=34 divde by 4 on both sides bcuz you want x by itself x=8.5

OpenStudy (anonymous):

Ohhh ok I get it ! Thank you soo much ! U make more sense then others.

OpenStudy (anonymous):

No problem :)